KIllerwolfdaone:

The equation above shows how temperature F, measured in degrees Fahrenheit, relates to a temperature C, measured in degrees Celsius. Based on the equation, which of the following must be true? A temperature increase of 1 degree Fahrenheit is equivalent to a temperature increase of 5 9 degree Celsius. A temperature increase of 1 degree Celsius is equivalent to a temperature increase of 1.8 degrees Fahrenheit. A temperature increase of 5 9 degree Fahrenheit is equivalent to a temperature increase of 1 degree Celsius. A) I only B) II only C) III only D) I and II only

KIllerwolfdaone:

nope

Honda:

this may help a little.

KIllerwolfdaone:

thx

AZ:

@Killerwolfdaone Do you still need help? I've answered this question on QuestionCove three times already lol

AZ:

Let's take a look at case 1 first We increase the temperature in Fahrenheit by 1, will the temperature in Celsius change by 5/9? let's see C=5/9(F−32) Let's first say that we have 32 Fahrenheit what would you get? C = 5/9 (32- 32) C = ?

AZ:

And after you calculate that, let's use 33 Fahrenheit. So that way we can test option 1. What happens when we increase the temperature by 1 F So we have C = 5/9 (33-32) C = ?? Is the difference between the first answer and this one 5/9? If so, that makes condition 1 true